# Single color virtual H&E staining with In-and-Out Net

### v1

@@ 2.1 Data collection
Twenty skin tissue samples were collected from sixteen patients undergoing Mohs micrographic surgery. Eight of these samples exhibited BCC and normal tissue, while the remaining twelve contained solely normal tissue. Each sample underwent frozen sectioning of 8 µm thickness at -22℃ with a microtome. After sectioning, a solution composed of acridine orange (AO) and Sulforhodamine 101 (SR101) (8mg AO and 5mg SR101 in 50ml 1:1 water:ethanol solution) was applied for 2 minutes on the sectioned skin samples, followed by 30-second rinse with a 1:1 water:ethanol solution. Afterward, a 30% aluminum chloride water solution was applied for 1 minute, followed by a rinse with pure ethanol. Aluminum chloride increases nucles contrast under RCM and is a chemical applied to stop bleeding during surgery.Scope et al. (2010); Flores et al. (2015) Following the staining process, confocal fluorescence microscopy mosaic images were obtained using Bruker Prairie View Ultima IV. Each scan covered a 450*450 µm field of view with 512*512 pixels. As illustrated in Fig. 1, each scanning procedure captured 5-15 different depth layers with a step size of 1 µm, resulting in an image stack. The excitation wavelengths were 488nm and 561nm. 488nm laser was used for reflectance mode and AO excitation, and 561nm was used for SR101 excitation. Three channels were employed for data collection: the reflectance channel for wavelengths <495nm, the AO channel for wavelengths 525±50nm, and the SR101 channel for wavelengths 605±70nm. The dataset comprises 705 image stacks within each channel. This research was approved by the Institutional Review Board at The University of Texas at Austin and the Seton Healthcare Family.
@@ 2.2 Data preprocessing
The image stacks represent different layers of the sample. In many scenarios, achieving a perfectly flat sample surface is challenging, making it difficult to ensure precise alignment of the laser tomography sectioning plane with the sample surface. Consequently, extracting surface information from image stacks becomes necessary. One effective strategy involves leveraging the stacked images as model inputs, allowing the network to discern surface details Lotfollahi et al. (2019); Li et al. (2021). However, in the future where virtual staining can be applied in vivo, patient wounds are unlikely to exhibit the same flatness as sectioned tissue, potentially resulting in each layer containing diminished information compared to sectioned tissue samples. Moreover, using image stacks as model inputs mandates consistency in the number of layers (the number of images within each stack), imposing limitations on the layer numbers. To address this challenge, we adopted the surface extraction algorithm proposed by Gao et al. Gao et al. (2023) to extract a single-layer image from each stack. This algorithm calculates the Fourier transformation along depth axis to find the rapid change pixels and determine the surface.
@@
Furthermore, an optical interference issue caused a bright dot to appear in all reflectance channel images (Fig. 1). The Bruker Prairie View Ultima IV system isn’t originally designed for reflectance mode. When we replace the dichroic with a beam splitter to enable both fluorescence and reflectance modes, the unintended artifact of reflected light within the optical path emerges within the reflectance channels. To address this, we acquired an additional image specifically capturing the interference dot, serving as a mask. Subsequently, we employed image inpainting techniques to fill in and correct the bright dot in the images. AO and SR101 channel images were transformed into H&E images as ground truth using the method proposed by Giacomelli et al. Giacomelli et al. (2016). During the training phase of our experiment, each image underwent random cropping, resulting in the creation of 10 distinct batches of 256x256 pixels. We collected 705 images from 15 patients (8,225,15,32,105,12,11,103,40,5,13,9,45,12,70 images from each patient, respectively). The variation in numerical data among patients came from the inconsistent imaging quality. As previously mentioned, our imaging system wasn’t initially designed for reflectance mode, resulting in significant variability in RCM image quality. We filtered out images based on four criteria: 1. Overexposed and underexposed images with substantial loss of information; 2. Excessive exposure disparities between RCM and fluorescence images; 3. Images lacking clear nuclei contrast under RCM due to poor aluminum chloride staining; 4. Outside region of interest images and corrupted images. Finally, we used 13 patients with 587 images as training set and 2 patients with 118 images as test set. After random cropping, we end up with 5870 images in training set and 1180 images in test set.
@@ 2.3 Model Architecture
The In-and-Out Net architecture is illustrated in Fig. 2. Fig. 2a depicts the overall architecture of the network, while Fig. 2b describes the inner loop and outer loop of the network. Additionally, Fig. 2c provides a detailed structure for both the generator and discriminator.
@@ 2.3.1 Overall architecture
Figure 2a illustrates the overall architecture of the In-and-Out Net. The network comprises two generators, three discriminators, and a trainable RGB concatenate layer. When dealing with H&E-stained skin tissue, hematoxylin predominantly stains nuclei in blue-purple, while eosin stains cytoplasm, connective tissue, and extracellular matrix components in pink-red. Our network aims to generate images stained separately with these two chemicals, referred to as the ”H channel” image and the ”E channel” image for hematoxylin and eosin staining, respectively. The two generators, $G_{H}$ and $G_{E}$, are responsible for generating the H and E channel images, each with a corresponding discriminator, $D_{H}$ and $D_{E}$. The RGB image is then generated from the H and E channel images using a trainable RGB concatenate layer, as shown in Eq. 1.
$\mathbf{I}_{RGB}(i,j,c)$ $=1-\mathbf{I}_{H}(i,j)\cdot\text{sigmoid}(W_{H,c})-\mathbf{I}_{E}(i,j)\cdot\text{sigmoid}(W_{E,c})$ (1)
@@
Here, $\mathbf{I}_{H}(i,j)$ and $\mathbf{I}_{E}(i,j)$ represent the pixel values of the output single-channel images from $G_{H}$ and $G_{E}$, respectively. $\mathbf{I}_{RGB}(i,j,c)$ is the pixel value of the output three-channel image from the RGB Concatenate Layer. $W_{H,c}$ and $W_{E,c}$ are trainable variables, where $c=0,1,2$ corresponds to $R,G,B$ channels, respectively. For the output image $\mathbf{I}_{RGB}(i,j,c)$, an additional discriminator $D_{out}$ is employed for supervision. Eq. 1 can also be written as:
@@ 2.3.2 Inner loop and outer loop
As depicted in Fig. 2a, $G_{H}$, $G_{E}$, $\mathbf{I}_{H}$, $\mathbf{I}_{E}$, and the RGB concatenate layer are enclosed within a square box, forming a larger generator denoted as $G_{out}$. Within $G_{out}$, referred to as the inner loop, the combination of $G_{H}$, $G_{E}$, $\mathbf{I}_{H}$, $\mathbf{I}_{E}$, and the RGB concatenate layer collectively functions as a single entity corresponding to $D_{out}$ in the outer loop.
During the training process, the inner loop and outer loop are trained separately. Fig. 2b illustrates the training of the inner loop, where the training process exclusively follows the red arrows. In this phase, $G_{H}$ and $G_{E}$ are trained with the assistance of $D_{H}$ and $D_{E}$ to generate $\mathbf{I}_{H}$ and $\mathbf{I}_{E}$. The inner loop emphasizes the training of structural information in the output images, excluding considerations of color information and the process of combining the structural images. Consequently, it serves as a more detailed structure-oriented loop.
@@ 2.3.3 Generator and discriminator architectures
The generator and discriminator architectures are shown in Fig. 2c. The generators, $G_{H}$ and $G_{E}$, share identical architectures. Both generators employ a U-Net structure based on the pix2pixGAN Generator framework. The U-Net architecture is known for its effectiveness in image-to-image translation tasks, allowing for the preservation of high-level structural features while generating realistic outputs. This design choice enhances the network’s ability to capture and reproduce intricate details in the H and E channel images.
@@ 2.4 Training and implementation
The optimization of the generator ($G$) and discriminator ($D$) is guided by the pix2pixGAN loss functionIsola et al. (2017):
@@
Here, $x$ represents the condition (RCM image), $z$ is the input image (RCM image), and $y$ is the ground truth image (H&E-stained image). In this context, when $x=z$, the generator takes a single RCM image as input. The generator aims to minimize the overall loss, while the discriminator seeks to maximize it. The first two terms originate from the standard conditional GAN (cGAN)Mirza and Osindero (2014) loss, and the third term introduces the $\mathcal{L}_{1}$ loss to encourage the generated image to closely resemble the ground truth image. The weight of the $\mathcal{L}_{1}$ loss is denoted as $\lambda_{0}$.
@@ 2.4.2 Inner and outer loop loss
$L_{\text{out}}$ =$&L_{\text{pix2pix}}(G_{\text{out}},D_{\text{out}})+\\ &\lambda_{1}\cdot\mathbb{E}_{x,y}[\|y-G_{H}(x)\|_{1}]+\lambda_{2}\cdot\mathbb{E}_{x,y}[\|y-G_{E}(x)\|_{1}]+\\ &\mathbb{E}_{x}[\log(1-D_{H}(x,G_{H}(x)))]+\mathbb{E}_{x}[\log(1-D_{E}(x,G_{E}(x)))]\end{aligned}$ (6)
@@
The first line in Eq. 6 represents the standard pix2pixGAN loss. Additionally, $\mathcal{L}_{1}$ loss terms for both $G_{H}$ and $G_{E}$ are included in the second line to preserve structural information during the outer loop. Moreover, $D_{H}$ and $D_{E}$ contribute to the third line, serving a role similar to the $\mathcal{L}_{1}$ loss of $G_{H}$ and $G_{E}$, preventing significant changes in structural information. However, $D_{H}$ and $D_{E}$ are not trained during the outer loop with their weights frozen. The function $f_{\text{RGB\_Concat}}$ is defined in Eq. 3.
@@ 2.4.4 Implementation
The training process follows a specific pattern, consisting of 10 epochs in the inner loop, followed by 10 epochs in the outer loop, and then another 10 epochs in the inner loop. This alternating training scheme continues until a total of 400 epochs are completed. Throughout the training, the step size remains consistently at $10^{-4}$, and the batch size remains fixed at 16 batches. The Adam optimizer is employed as the trainer for both the inner and outer loops. The code is written in the Python programming language with PyTorch and executed on an Nvidia RTX 3070 GPU for 10 hours.
@@ 2.5.1 Model comparison
We compare our In-and-Out Net with four existing models: pix2pixGANIsola et al. (2017), pix2pixHDWang et al. (2018), VSGD-NetLiu et al. (2023), and RestainNetZhao et al. (2022). pix2pixGAN serves as the foundational structure for our network, while pix2pixHD builds upon pix2pixGAN by incorporating additional blocks and introducing object segmentation. VSGD-Net is specifically designed for transforming H&E-stained images into Sox10 stained images, focusing on the specification of melanocytes. In our adaptation, we replace melanocytes with RCM nuclei to enhance nuclear staining. RestainNet addresses the stain normalization task, transforming one H&E-stained image into another with a different color style. Its architecture closely resembles ours. Instead of initially training the input image to H and E channel images, RestainNet first trains the input image to an H&E-stained image. Subsequently, it breaks down this H&E-stained image into H and E channels, feeding these two channel images back into the loss function. To assess and compare the performance of these models, we utilize metrics such as Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), Feature Similarity Index (FSIM), and Multi-Scale Structural Similarity Index (MS-SSIM).
@@ 2.5.3 Training approach comparison
The objective of alternating training between the inner and outer loops is to stabilize structural details (inner loop) and subsequently make fine adjustments (outer loop) to color and channel combinations. To delve deeper into this training method, including its sensitivity to epoch numbers, we conducted a detailed study. To effectively compare the training approaches, monitoring the loss changes during training is essential. However, utilizing the overall loss function from Eq. 8 alone is inaccurate, given that inner and outer loops are dominated by different loss functions. To address this, we used the following function, denoted as $eval\_loss$:

### v2

@@ 2.1 Data collection
Twenty skin tissue samples were collected from sixteen patients undergoing Mohs micrographic surgery. Eight of these samples exhibited BCC and normal tissue, while the remaining twelve contained solely normal tissue. Each sample underwent frozen sectioning of 8 µm thickness at -22℃ with a microtome. After sectioning, a solution composed of acridine orange (AO) and Sulforhodamine 101 (SR101) (8mg AO and 5mg SR101 in 50ml 1:1 water:ethanol solution) was applied for 2 minutes on the sectioned skin samples, followed by 30-second rinse with a 1:1 water:ethanol solution. Afterward, a 30% aluminum chloride water solution was applied for 1 minute, followed by a rinse with pure ethanol. Aluminum chloride increases nucles contrast under RCM and is a chemical applied to stop bleeding during surgery.(Scope et al. (2010); Flores et al. (2015) Following the staining process, confocal fluorescence microscopy mosaic images were obtained using Bruker Prairie View Ultima IV. Each scan covered a 450*450 µm field of view with 512*512 pixels. As illustrated in Figure 1, each scanning procedure captured 5-15 different depth layers with a step size of 1 µm, resulting in an image stack. The excitation wavelengths were 488nm and 561nm. 488nm laser was used for reflectance mode and AO excitation, and 561nm was used for SR101 excitation. Three channels were employed for data collection: the reflectance channel for wavelengths <495nm, the AO channel for wavelengths 525±50nm, and the SR101 channel for wavelengths 605±70nm. The dataset comprises 705 image stacks within each channel. This research was approved by the Institutional Review Board at The University of Texas at Austin and the Seton Healthcare Family.
@@ 2.2 Data preprocessing
The image stacks represent different layers of the sample. In many scenarios, achieving a perfectly flat sample surface is challenging, making it difficult to ensure precise alignment of the laser tomography sectioning plane with the sample surface. Consequently, extracting surface information from image stacks becomes necessary. One effective strategy involves leveraging the stacked images as model inputs, allowing the network to discern surface details (Lotfollahi et al. (2019); Li et al. (2021)). However, in the future where digital staining can be applied in vivo, patient wounds are unlikely to exhibit the same flatness as sectioned tissue, potentially resulting in each layer containing diminished information compared to sectioned tissue samples. Moreover, using image stacks as model inputs mandates consistency in the number of layers (the number of images within each stack), imposing limitations on the layer numbers. To address this challenge, we adopted the surface extraction algorithm proposed by Gao et al. (Gao et al. (2023)) to extract a single-layer image from each stack. This algorithm calculates the Fourier transformation along depth axis to find the rapid change pixels and determine the surface.
@@
Furthermore, an optical interference issue caused a bright dot to appear in all reflectance channel images (Figure 1). The Bruker Prairie View Ultima IV system isn’t originally designed for reflectance mode. When we replace the dichroic with a beam splitter to enable both fluorescence and reflectance modes, the unintended artifact of reflected light within the optical path emerges within the reflectance channels. To address this, we acquired an additional image specifically capturing the interference dot, serving as a mask. Subsequently, we employed image inpainting techniques to fill in and correct the bright dot in the images. AO and SR101 channel images were transformed into H&E images as ground truth using the method proposed by Giacomelli et al. (Giacomelli et al. (2016)). During the training phase of our experiment, each image underwent random cropping, resulting in the creation of 10 distinct batches of 256x256 pixels. We collected 705 images from 15 patients (8,225,15,32,105,12,11,103,40,5,13,9,45,12,70 images from each patient, respectively). The variation in numerical data among patients came from the inconsistent imaging quality. As previously mentioned, our imaging system wasn’t initially designed for reflectance mode, resulting in significant variability in RCM image quality. We filtered out images based on four criteria: 1. Overexposed and underexposed images with substantial loss of information; 2. Excessive exposure disparities between RCM and fluorescence images; 3. Images lacking clear nuclei contrast under RCM due to poor aluminum chloride staining; 4. Outside region of interest images and corrupted images. Finally, we used 13 patients with 587 images as training set and 2 patients with 118 images as test set. After random cropping, we end up with 5870 images in training set and 1180 images in test set.
@@ 2.3 Model Architecture
The In-and-Out Net architecture is illustrated in Figure 2. Figure 2a depicts the overall architecture of the network, while Figure 2b describes the inner loop and outer loop of the network. Additionally, Figure 2c provides a detailed structure for both the generator and discriminator.
@@ 2.3.1 Overall architecture
Figure 2a illustrates the overall architecture of the In-and-Out Net. The network comprises two generators, three discriminators, and a trainable RGB concatenate layer. When dealing with H&E-stained skin tissue, hematoxylin predominantly stains nuclei in blue-purple, while eosin stains cytoplasm, connective tissue, and extracellular matrix components in pink-red. Our network aims to generate images stained separately with these two chemicals, referred to as the ”H channel” image and the ”E channel” image for hematoxylin and eosin staining, respectively. The two generators, $G_{H}$ and $G_{E}$, are responsible for generating the H and E channel images, each with a corresponding discriminator, $D_{H}$ and $D_{E}$. The RGB image is then generated from the H and E channel images using a trainable RGB concatenate layer, as shown in Equation 1.
$\mathbf{I}_{RGB}(i,j,c)$ $=1-\mathbf{I}_{H}(i,j)\cdot\text{sigmoid}(W_{H,c}-\mathbf{I}_{E}(i,j)\cdot\text{sigmoid}(W_{E,c}$ (1)
@@
Here, $\mathbf{I}_{H}(i,j)$ and $\mathbf{I}_{E}(i,j)$ represent the pixel values of the output single-channel images from $G_{H}$ and $G_{E}$, respectively. $\mathbf{I}_{RGB}(i,j,c)$ is the pixel value of the output three-channel image from the RGB Concatenate Layer. $W_{H,c}$ and $W_{E,c}$ are trainable variables, where $c=0,1,2$ corresponds to $R,G,B$ channels, respectively. For the output image $\mathbf{I}_{RGB}(i,j,c)$, an additional discriminator $D_{out}$ is employed for supervision. Equation 1 can also be written as:
@@ 2.3.2 Inner loop and outer loop
As depicted in Figure 2a, $G_{H}$, $G_{E}$, $\mathbf{I}_{H}$, $\mathbf{I}_{E}$, and the RGB concatenate layer are enclosed within a square box, forming a larger generator denoted as $G_{out}$. Within $G_{out}$, referred to as the inner loop, the combination of $G_{H}$, $G_{E}$, $\mathbf{I}_{H}$, $\mathbf{I}_{E}$, and the RGB concatenate layer collectively functions as a single entity corresponding to $D_{out}$ in the outer loop.
During the training process, the inner loop and outer loop are trained separately. Figure 2b illustrates the training of the inner loop, where the training process exclusively follows the red arrows. In this phase, $G_{H}$ and $G_{E}$ are trained with the assistance of $D_{H}$ and $D_{E}$ to generate $\mathbf{I}_{H}$ and $\mathbf{I}_{E}$. The inner loop emphasizes the training of structural information in the output images, excluding considerations of color information and the process of combining the structural images. Consequently, it serves as a more detailed structure-oriented loop.
@@ 2.3.3 Generator and discriminator architectures
The generator and discriminator architectures are shown in Figure 2c. The generators, $G_{H}$ and $G_{E}$, share identical architectures. Both generators employ a U-Net structure based on the pix2pixGAN Generator framework. The U-Net architecture is known for its effectiveness in image-to-image translation tasks, allowing for the preservation of high-level structural features while generating realistic outputs. This design choice enhances the network’s ability to capture and reproduce intricate details in the H and E channel images.
@@
The layer settings, parameter numbers and training settings are present in Table 1.
Table 1: Layer settings, activation functions, and optimization methods for In-and-Out Net (Generator and Discriminator)
Layer Type (Generator) Output Shape Kernel Size Stride Activation Function
Input Layer (3, 256, 256) - - -
Initial Convolution (Downsample) (64, 128, 128) 4x4 2 LeakyReLU
Convolutional Layer 1 (Downsampling) (128, 64, 64) 4x4 2 LeakyReLU
Convolutional Layer 2 (Downsampling) (256, 32, 32) 4x4 2 LeakyReLU
Convolutional Layer 3 (Downsampling) (512, 16, 16) 4x4 2 LeakyReLU
Bottleneck (512, 8, 8) 4x4 2 ReLU
Transpose Conv Layer 1 (Upsampling) (512, 16, 16) 4x4 2 ReLU (with Dropout)
Transpose Conv Layer 2 (Upsampling) (512, 32, 32) 4x4 2 ReLU (with Dropout)
Transpose Conv Layer 3 (Upsampling) (256, 64, 64) 4x4 2 ReLU
Final Transpose Conv Layer (3, 256, 256) 4x4 2 Tanh
Total Parameters (Generator) 108,828,940 (In-and-Out Net)
Layer Type (Discriminator) Output Shape Kernel Size Stride Activation Function
Input Layer (3, 256, 256) - - -
Convolutional Layer 1 (64, 128, 128) 4x4 2 LeakyReLU
Convolutional Layer 2 (128, 64, 64) 4x4 2 LeakyReLU
Convolutional Layer 3 (256, 32, 32) 4x4 2 LeakyReLU
Convolutional Layer 4 (512, 16, 16) 4x4 2 LeakyReLU
Convolutional Layer 5 (512, 8, 8) 4x4 2 LeakyReLU
Final Convolutional Layer (1, 1, 1) 4x4 1 Sigmoid
Total Parameters (Discriminator) 2,765,121
Optimization Method Setting
Optimizer Adam
Learning Rate 0.0002
Beta1 (Momentum term) 0.5
Weight Decay 0 (None)
Loss Function GAN Loss (pix2pix framework)
Epochs 400
Batch Size 16
@@ 2.4 Training and implementation
The optimization of the generator ($G$) and discriminator ($D$) is guided by the pix2pixGAN loss function(Isola et al. (2017)):
@@
Here, $x$ represents the condition (RCM image), $z$ is the input image (RCM image), and $y$ is the ground truth image (H&E-stained image). In this context, when $x=z$, the generator takes a single RCM image as input. The generator aims to minimize the overall loss, while the discriminator seeks to maximize it. The first two terms originate from the standard conditional GAN (cGAN)(Mirza and Osindero (2014)) loss, and the third term introduces the $\mathcal{L}_{1}$ loss to encourage the generated image to closely resemble the ground truth image. The weight of the $\mathcal{L}_{1}$ loss is denoted as $\lambda_{0}$.
@@ 2.4.2 Inner and outer loop loss
$L_{\text{out}}$ =$&L_{\text{pix2pix}}(G_{\text{out}},D_{\text{out}}+\\ &\lambda_{1}\cdot\mathbb{E}_{x,y}[\|y-G_{H}(x)\|_{1}]+\lambda_{2}\cdot\mathbb{E}_{x,y}[\|y-G_{E}(x)\|_{1}]+\\ &\mathbb{E}_{x}[\log(1-D_{H}(x,G_{H}(x)))]+\mathbb{E}_{x}[\log(1-D_{E}(x,G_{E}(x)))]\end{aligned}$ (6)
@@
The first line in Equation 6 represents the standard pix2pixGAN loss. Additionally, $\mathcal{L}_{1}$ loss terms for both $G_{H}$ and $G_{E}$ are included in the second line to preserve structural information during the outer loop. Moreover, $D_{H}$ and $D_{E}$ contribute to the third line, serving a role similar to the $\mathcal{L}_{1}$ loss of $G_{H}$ and $G_{E}$, preventing significant changes in structural information. However, $D_{H}$ and $D_{E}$ are not trained during the outer loop with their weights frozen. The function $f_{\text{RGB\_Concat}}$ is defined in Equation 3.
@@ 2.4.4 Implementation
The training process follows a specific pattern, consisting of 10 epochs in the inner loop, followed by 10 epochs in the outer loop, and then another 10 epochs in the inner loop. This alternating training scheme continues until a total of 400 epochs are completed. Throughout the training, the step size remains consistently at $10^{-4}$, and the batch size remains fixed at 16 batches. The Adam optimizer is employed as the trainer for both the inner and outer loops. The code is written in the Python programming language with PyTorch and executed on an Nvidia RTX 3060 GPU for 16 hours.
@@ 2.5.1 Model comparison
We compare our In-and-Out Net with four existing models: pix2pixGAN(Isola et al. (2017)), pix2pixHD(Wang et al. (2018)), VSGD-Net(Liu et al. (2023)), and RestainNet(Zhao et al. (2022)). pix2pixGAN serves as the foundational structure for our network, while pix2pixHD builds upon pix2pixGAN by incorporating additional blocks and introducing object segmentation. VSGD-Net is specifically designed for transforming H&E-stained images into Sox10 stained images, focusing on the specification of melanocytes. In our adaptation, we replace melanocytes with RCM nuclei to enhance nuclear staining. RestainNet addresses the stain normalization task, transforming one H&E-stained image into another with a different color style. Its architecture closely resembles ours. Instead of initially training the input image to H and E channel images, RestainNet first trains the input image to an H&E-stained image. Subsequently, it breaks down this H&E-stained image into H and E channels, feeding these two channel images back into the loss function. To assess and compare the performance of these models, we utilize metrics such as Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), Feature Similarity Index (FSIM), and Multi-Scale Structural Similarity Index (MS-SSIM).
@@ 2.5.3 Training approach comparison
The objective of alternating training between the inner and outer loops is to stabilize structural details (inner loop) and subsequently make fine adjustments (outer loop) to color and channel combinations. To delve deeper into this training method, including its sensitivity to epoch numbers, we conducted a detailed study. To effectively compare the training approaches, monitoring the loss changes during training is essential. However, utilizing the overall loss function from Equation 8 alone is inaccurate, given that inner and outer loops are dominated by different loss functions. To address this, we used the following function, denoted as $eval\_loss$: